OpenStudy (anonymous):

simplify the rational expression: x+4 --------- x^2+5x+4

jhonyy9 (jhonyy9):

how can you factotizing x^2 +5x +4 =0 ?

jhonyy9 (jhonyy9):

do you know it ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

:(

jhonyy9 (jhonyy9):

so one little help x*x +x +4x +4 =0 than now what can you factorizing out from the first two terms and from 3rd and 4rd ?

jhonyy9 (jhonyy9):

x(x+1) +4(x+1)=0 yes ? than now you continue and factoriz out the (x+1) and will get (x+1)(x+4)=0

jhonyy9 (jhonyy9):

so using this you will can rewriting your exercise how ?

jhonyy9 (jhonyy9):

x+4 ---------- = ... (x+1)(x+4) hope so much that you will can finich this ,yes ?

OpenStudy (anonymous):

wouldnt the x+4 on the top and bottom cancel out?

jhonyy9 (jhonyy9):

yes simplifie it easy

OpenStudy (anonymous):

so its just x+1?

jhonyy9 (jhonyy9):

why just x+1 ?

OpenStudy (anonymous):

idk...im asking

jhonyy9 (jhonyy9):

so if you cancel out the (x+4) from top and bottom ,so from numerator and denominator what will reamining there ?

OpenStudy (anonymous):

nothing on the top and the x+1 on the bottom?

jhonyy9 (jhonyy9):

when you simplifie 2/4 = how many ? this is just an example for you can understanding it why not is right your answer

OpenStudy (anonymous):

1/2...

jhonyy9 (jhonyy9):

yes right so than in case of your exercise here npw what will reamining ?

OpenStudy (anonymous):

i dont know..i dont understand

jhonyy9 (jhonyy9):

why ? if you simplifie 2/4 this is equal 1/2 so than if you simplifie x+4 ----------- = ---------------- (x+1)(x+4) so how many will reamaning on the top ? nothing is not possible

OpenStudy (anonymous):

i thought that x+4 cancelled out on the top and bottom???

jhonyy9 (jhonyy9):

yes is right but in my example when there are 2/4 so there canceled out 2 and top and bottom too but your answer was that resulted here in this case 1/2 so than why ?

jhonyy9 (jhonyy9):

why 1/2 and not just 2 ?

OpenStudy (anonymous):

so it would be 1 ---- (x+1)

jhonyy9 (jhonyy9):

yes sure this is the right answer but do you understand it now ?

OpenStudy (anonymous):

i understand this one lol...i have a lot of problems to do still but i have class at 1130 ):

OpenStudy (anonymous):

thank you (:

jhonyy9 (jhonyy9):

was my pleasure good luck bye